# A Geometric Tension Dynamics Model of Epithelial Convergent Extension

### v2

@@ 4.3 Cell shape elasticity
As explained in the main text, in our model, tension dynamics translate into tissue dynamics via force balance which we implement by minimizing the energy function Eq. (1). Here, we explain our choice of the term representing cell-shape elasticity, $E_{C}$. We use an elastic energy based on the cell shape tensor
@@
with bulk modulus $\lambda$ and shear modulus $\mu$. The reference shape $S_{0}$ controls the isogonal mode. An isotropic $S_{0}\propto\mathrm{Id}$ favours equilateral hexagons. We used a shear/bulk ration of $\mu/\lambda=1$ for all “active” cells in the simulations. Because of the separation of scales between cortical tensions and elastic energy baked into the model, the absolute values of $\lambda,\mu$ are irrelevant.
Strikingly, when the cell array is a regular lattice (see Fig. 1G and single-quartet simulations shown in the companion paper [17]), the “shape strain” $S_{\mathcal{C}}-S_{0}$ can always be relaxed to 0 by choice of the edge lengths, and the energy $E_{\mathcal{C}}=0$ throughout. This can already be seen from a degree-of-freedom count (three $\ell_{i}$ for the three independent components of $S_{\mathcal{C}}-S_{0}$). The elastic energy therefore acts only on the isogonal modes (i.e. $\partial_{\phi_{i}}E_{\mathcal{C}}|_{\ell_{i}=\mathrm{minimizers}}=0$) and there is no energy barrier for intercalations. Consequently, there is no need for noise to drive intercalations in our model. By contrast, for the widely-used area-perimeter elastic energy $E=(A-A_{0})^{2}+(P-P_{0})^{2}$ (where $A,P$ are the cell area and perimeter, and $A_{0},P_{0}$ their target values) [29], there exists an energy barrier, and the inner interface $\ell_{0}$ only collapses when $\beta\to\pi$ (see Fig. S14B). Note that the area-perimeter energy is a special case because of the geometric incompatibility of area and perimeter constraints when $P_{0}/\sqrt{A_{0}}<3.72$. For $P_{0}/\sqrt{A_{0}}>3.72$ cell shape is underdetermined (floppy) and the shear modulus vanishes. Recall that we have shown that a finite shear modulus is required to translate tension dynamics into tissue-level convergent extension (see Fig. 8G). The area-perimeter energy in the fluid regime is therefore not suitable to model convergent extension driven by internal tension dynamics.
Combining area elasticity with shear elasticity based on the shape tensor, $E=(A-A_{0})^{2}+\mu\mathrm{Tr}[(S_{\mathcal{C}}-S_{0})^{2}]$, (or perimeter elasticity with cell shape bulk elasticity) leads to similar results as Eq. (3). Note also that because of the degree-of-freedom count, the system is under-specified if the shear modulus is 0, foreshadowing the fact that without shear modulus, no convergence-extension takes place. Indeed, in the case of the symmetric lattice, the degrees of freedom are the three vertex angles, determined by the tensions, and the three interface lengths, independent of the angles. Fixing the cell area still leaves freedom to change the interface lengths. More generally, we have shown above and illustrated in Fig. 1 that isogonal modes allow for area-preserving shear transformations.
@@ 4.4 Tissue scale simulation
Using the components of the single-quartet simulation, we now describe a simulation of an arbitrary, disordered tissue patch. The key assumption are again adiabatic force balance and cortical tension dominance: the tissue geometry is found by solving for force balance, by first finding the angles from the cortical tensions and then the isogonal mode by minimizing cell elastic energy. The key quantity is the tension triangulation whose dynamics determines the tissue dynamics.
@@
We model a piece of epithelial tissue by a polygonal tiling of the plane, polygons corresponding to cells. Two cells are neighbors if they are share an edge (this implies that the cell adjacency graph is a triangulation).
@@
  • Faces: primal coordinates, the $x,y$ coordinates of the vertex in the cell tessellation corresponding to the face, denoted $\mathbf{r}_{ijk}$.
@@
The mesh is implemented in object oriented python, using a half-edge mesh data structure [74]. A half-edge mesh consists of vertices, faces, and, for each edge, two oriented half-edges. A half-edges stores a reference to the next half-edge on the counter-clockwise oriented face it belongs to, and a reference to its twin, the half-edge with opposite orientation on the adjacent face. This structure defines an orientation for every triangular face and makes mesh traversal (e.g. get all vertices of a cell) and mesh modification (e.g. intercalations) very convenient.
@@
Due to the assumption of adiabatic force balance, the tissue geometry is determined by the instantaneous tensions. The dynamics is therefore determined by the change in the tension triangulation (due to positive feedback). In detail, the time evolution in our simulation is done in three steps:
@@
    Triangulation flattening: Optimize the dual vertex positions so that the lengths of each half-edge is as close as possible to its internally stored tension. This ensures global tension balance.
@@
    Cell shape optimization: Optimize the primal vertex positions to minimize the shape-tensor based elastic energy, while constraining the angles by the dual triangulation.
@@
The next sections describe them in detail. Fig. S15 provides a summary of the simulation workflow. The most complicated and time-consuming of them is the cell shape optimization. Overall, a simulation of convergent extension of $\sim 1000$ cells takes approx. 15 minutes on a current-generation laptop.
Note that the workflow of our simulations differs somewhat from the more common used dissipative dynamics, in which cell geometry is updated dynamically according to the ODE $\gamma\partial_{t}\mathbf{r}_{ijk}=-\frac{\partial E}{\partial\mathbf{r}_{ijk}}$. The dual separation of scales (fast relaxation and dominance of cortical tensions), mean that a direct implementation of dissipative dynamics is numerically challenging. The elastic energy minimization can be seen as a relaxational Euler step with an adaptive step size.
Relaxation via fast dissipative dynamics is a viable approach that could indeed be used as an alternative numerical relaxation method. However, as the theory of fast variable elimination in the context of multiscale dynamics is well established [75], we do not expect a different scheme to affect our results to an important degree. The key “ingredient” of our model is not how exactly the elastic energy is relaxed, but how active tensions are regulated.
Time stepping Choose time-step size $dt$. For $i=1,...,N_{\mathrm{steps}}$ do: 1. Update intrinsic tensions (positive feedback Euler step): $\tilde{T}_{ij}\mapsto\tilde{T}_{ij}+f(\tilde{T}_{ij},...)dt$. $f$ is the feedback RHS. 2. Update tension triangulation vertices (least-squares fit to intrinsic tensions): $\mathbf{t}_{i}\mapsto\mathrm{argmin}\frac{1}{2N_{\mathrm{edges}}}\sum_{ij}(|\mathbf{t}_{i}-\mathbf{t}_{j}|-\tilde{T}_{ij})^{2}+\frac{1}{2N_{\mathrm{tri}}}\ \sum_{ijk}\mathrm{Pen}(A_{ijk},A_{0})$ 3. Update intrinsic tensions (flattness feedback Euler step): $\tilde{T}_{ij}\mapsto\tilde{T}_{ij}-\tau_{\mathrm{balance}}^{-1}(|\mathbf{t}_{i}-\mathbf{t}_{j}|-\tilde{T}_{ij})dt$ 4. Update cell vertices by angle-constrained energy minimization: $\mathbf{r}_{ijk}\mapsto\mathrm{argmin}\frac{1}{2N_{\mathrm{tri}}}\sum_{ij}T_{ij}\overline{\ell}(1-\hat{n}_{ij}\cdot\mathbf{r}_{ij})+\frac{\varepsilon}{2N_{\mathrm{cells}}}\sum_{i}E_{\mathcal{C},i}(S_{i})$ 5. Update topology: flip all edges of length $\ell_{ij}<\ell_{\mathrm{min}}$, update passive tensions $T_{p,ij}$ and intrinsic tensions $\tilde{T}_{ij}$ in flipped edges. 6. Log: save topology, tension vertices, cell vertices, and T1-transitions.
Figure S15: Simulation work flow showing a brief summary of the simulation code.
@@
The cutoff parameter $k_{\text{cutoff}}$ stops runaway positive feedback at values $T>1/k_{\text{cutoff}}$. It is not strictly necessary since runaway feedback is cut short by intercalations, but helpful for long-time simulations, were rarely, run-away tension cable configurations can occur which can blow up numerically. We chose $1/k_{\text{cutoff}}=5\bar{T}$ where $\bar{T}$ is the overall tension scale, so that it has almost no effect in normal tension configurations. We chose the feedback parameter $n=4$ and the relaxation rate $\tau_{p}=\tau_{T}/4$, the same used in the single-quartet and single-triangle simulations.
So far, the time evolution of tensions is purely autonomous (i.e. each tension triangle is independent from all others) and completely analogous to the single-quartet simulation. In the single-quartet simulation, all tension triangles are equal by symmetry and periodicity, and thus tile the plane as a triangulation by construction. In general, this is no longer guaranteed. Even if one starts with a flat tension triangulation, the autonomous time evolution Eq. (S25) will result in a set of tension triangles that cannot fit together to a flat triangulation. The triangulation will develop curvature [12], i.e. “crumple” out of plane. This means that we have to add an additional ingredient to enforce triangulation planarity.
We enforce planarity as follows. The set of all balanced tension triangulations is parameterized by the dual vertex positions $\mathbf{t}_{i}$. We therefore seek the balanced tension configuration that approximates the intrinsic tensions as closely as possible by minimizing the following elastic energy:
@@
Since the simulation is based on half-edges, the per-edge intrinsic tension $\tilde{T}_{ij}$ is the average of the tensions of the two half-edges (which can be though of as the two cellular cortices coupled in a cell-cell interface). The second term, $\mathrm{Pen}(A,A_{0})$ is a penalty term on the triangle areas $A$ (calculated from the vertex positions $\mathbf{t}_{i}$). It fixes the overall tension scale (which is arbitrary due to force balance) and ensures the triangulation is well behaved numerically. It contains two terms:
@@
In the simulation, we first carry out an explicit Euler step for the intrinsic tensions, and then “flatten the triangulation”, i.e. find the dual vertex positions by minimizing Eq. (S26). This determines the extrinsic tensions $T_{ij}=||\mathbf{t}_{i}-\mathbf{t}_{j}||$. Finally, we relax the intrinsic tensions to force balance with rate $\tau_{\mathrm{balance}}$,
@@
In order to ensure numerical stability even for large $\tau_{T}$, we use the exact solution of Eq. (S28) numerically, instead of carrying out an explicit Euler step.
The triangulation-flattening prescription can be thought of as modelling, indirectly, known feedback loops which lead to a convergence to balanced tension, such as strain-rate feedback [13], which has been demonstrated experimentally [14]. The rate $\tau_{\mathrm{balance}}^{-1}$ determines the speed of these feedback loops. The influence of the value of $\tau_{\mathrm{balance}}^{-1}$ is studied in Fig. S8. In principle, strain rate feedback requires movement of cell vertices to sense tension imbalance and then bring tensions back to a balanced state. We neglect this vertex motion, since we assume balance is restored rapidly. Directly modelling of strain rate feedback requires going beyond the separation between tension- and real-space-dynamics that is the basis of our model here and is therefore beyond the scope of this paper. Further, the triangulation-flattening prescription is the most parsimonious generalization of the local tension feedback dynamics employed in the single-quartet simulation to a disordered tissue patch.
The complete tension dynamics that is implemented in the simulation therefore comprises three terms in addition to the dynamics shown in the main text Eq. (5): (1) Numerical cutoff of positive feedback at very high tensions (2) Passive tension dynamics post T1 (3) Relaxation of tensions to force equilibrium:
@@
Above, we have described the dynamics of the tension triangulation. After an Euler step to update the tensions has been taken, we calculate the real space cell tessellation (i.e. the positions of the cell vertices) by minimizing the cell-shape elastic energy while constraining the angles by the tension triangulation. In the case of a single quartet of identical cells, we could explicitly parameterize the cell shape by interface lengths and angles, and optimize only with respect to the former. In the disordered case, it is more convenient to enforce the angle constraint via a (strong) energy penalty. It would also possible to instead directly optimize the per-cell isogonal potentials $\Theta_{i}$ with respect to the Voronoi reference, obviating the need for a penalty.
The elastic energy to minimize therefore has two terms – the cell elastic energy Eq. (3) calculated from cell shape tensor $S_{\mathcal{C}}$ Eq. (4.3) and reference shape $S_{0}$, and the angle penalty, which corresponds to the cortical tension energy term:
@@
Here, $\hat{n}_{ij}$ is the unit normal of the tension edge, $\hat{\mathbf{r}}_{ij}$ the unit tangent vector of the primal edge, and $\overline{\ell}$ is the average interface length, to ensure the term has the correct dimensions of energy. The small parameter $\varepsilon$ (we use $\varepsilon\sim 10^{-3}\lambda^{-1}$) represents the fact that the cell shape energy is assumed to be much weaker than the cortical tensions. If desired, one can replace or modify the cell shape elastic energy $E_{\mathcal{C}}$ in Eq. (S31) by another energy, e.g. the common area-perimeter elasticity.
In Eq. (S31), we use a different term than $T_{ij}\ell_{ij}$ to represent angle constraint due to cortical tensions. When determining the vertex positions $\mathbf{r}_{ijk}$, we minimize the cell shape elastic energy under the constraint of fixed angles, as given by the tension triangulation (so in a sense, we minimize the energy not with respect to the complete vertex positions, but with respect to the isogonal modes only). The tensions are externally imposed parameters and not functions of the vertex positions, like in a spring network. Eq. (S31) does not represent the physical energy, but merely the way we numerically enforce the angle constraints.
We use the alternative energy term $T_{ij}(1-\hat{n}_{ij}\cdot\hat{\mathbf{r}}_{ij})$ for two numerical reasons. (1) compared to $T_{ij}\ell_{ij}$, it does not require a pressure term $p\sum_{i}A_{i}$ to avoid global constriction; otherwise, we would have to solve a discrete Poisson equation to determine $p$. (2) it avoids potential convergence issues when minimizing the elastic energies in the presence of very short junctions, which occur before and after T1-transitions. However, if the overall area is fixed, both forms of the energy lead to the same family of ground states, namely states where the vertex angles are complementary to those of the tension triangulation. The energy term $T_{ij}(1-\hat{n}_{ij}\cdot\hat{\mathbf{r}}_{ij})$ was explicitly constructed to ensure this. Away from the ground state, the two energies differ; however, this is immaterial since we work in the regime where the tension constraints are always satisfied exactly.
We can also rewrite
@@
After carrying out a local Euler step and updating the dual vertex positions by “triangulation flattening”, we find the primal vertex positions by minimizing the elastic energy Eq. (S31).
@@
After each time-step we check whether any edge has collapsed. We call an edge collapsed if its length is $<10\%$ of the overall length scale (mean edge length at simulation start). If this occurs, we carry out an intercalation (flip the corresponding edge in the triangulation) and set the active and passive tensions on the new edge as described above.
Edges that underwent a T1 transition in the last 20 simulation time steps (corresponding to 2.5 minutes) are excluded from edge collapse so as to prevent the immediate re-collapse of newly formed junctions. Instead of this hard-coded criterion, one could also calculate the strain rate of junctions of short length and exclude extending junctions, as shown by the time traces of junction length in Fig. 8B.
@@
The simulation requires solving two optimization problems, one for the triangulation Eq. (S26) and one for the real-space shape Eq. (S31). To solve an optimization problem efficiently, it is crucial to know the gradient of the objective function. We use the python library JAX [76] to automatically differentiate our energy functions, allowing easy, rapid and bug-free calculation of the Jacobians. We further use JAX’s just-in-time functionality to accelerate our python code. Gradient-based optimization requires the objective function to be at least once differentiable. To ensure this, we mollify functions as required; for example, the Euclidean norm $\sqrt{x^{2}+y^{2}}\rightarrow\sqrt{x^{2}+y^{2}+\varepsilon^{2}},\>\varepsilon\sim 10^{-3}$ to ensure differentiability when computing the lengths of very short edges.
@@
We consider two types of initial conditions, using either a regular lattice or a random point process to choose the tension vertices. In the former case, we chose rectangular patches of a regular triangular lattice, with i.i.d. Gaussian noise added to the vertex positions. The amplitude of the noise was chosen so that $\mathrm{std}(T_{ij})=0.1\>\mathrm{mean}(T_{ij})$. The influence of the Gaussian noise standard deviation is shown in Fig. S8. A triangular lattice corresponds to a hexagonal lattice of cells. Indeed, locally, the Drosophila embryo at the onset of GBE looks like it is composed out of (noisy) hexagonal lattice patches, as shown in Fig. S1. This lattice is subjected to a small initial strain to create an a initial tension anisotropy, as described in the main text.
To investigate the effect of disorder, we also create initial conditions using a hard disk process. This randomly samples 2d-points (the eventual vertices in tension space) under the constraint that no two points are closer than $2r$ where $r$ is the “hard disk” radius. Instead of $r$, we use the normalized hard disk packing fraction $\rho\in[0,1]$, which is maximal for a hexagonal lattice. At $\rho\sim 0.72$, the hard disk process undergoes a phase transition that destroys the lattice order. The limit $\rho=0$ corresponds to a Poisson point process. To sample from a hard disk point process, we use the event-chain Monte Carlo algorithm of Ref. [31], with code based on that provided with Ref. [32]. Note that with other algorithms such as sequential deposition or standard Monte Carlo it is either impossible or prohibitively slow to sample at arbitrary values of $\rho$.
To create a tension triangulation, we then compute the Delaunay triangulation of this point set (we further refine the initial condition by removing the edge of this tension triangulation, where the Delaunay algorithm can create very long tension edges). This triangulation is stretched uniformly to set up the initial anisotropy of tension. Cell numbers and initial aspect ratios were also chosen identically in the simulations that make up the phase diagram of Fig. 3.

### v3

@@ 5.3 Cell shape elasticity
As explained in the main text, in our model, tension dynamics translate into tissue dynamics via force balance which we implement by minimizing the energy function (1). Here, we explain our choice of the term representing cell-shape elasticity, $E_{C}$. We use an elastic energy based on the cell shape tensor
@@
with bulk modulus $\lambda$ and shear modulus $\mu$. The reference shape $S_{0}$ controls the isogonal mode. An isotropic $S_{0}\propto\mathrm{Id}$ favours equilateral hexagons. We used a shear/bulk ratio of $\mu/\lambda=1$ for all “active” cells in the simulations. Because of the separation of scales between cortical tensions and elastic energy baked into the model, the absolute values of $\lambda,\mu$ are irrelevant.
Strikingly, when the cell array is a regular lattice (see Fig. 1G and single-quartet simulations shown in the companion paper [Brauns.etal2024elife]), the “shape strain” $S_{\mathcal{C}}-S_{0}$ can always be relaxed to 0 by choice of the edge lengths, and the energy $E_{\mathcal{C}}=0$ throughout. This can already be seen from a degree-of-freedom count (three $\ell_{i}$ for the three independent components of $S_{\mathcal{C}}-S_{0}$). The elastic energy therefore acts only on the isogonal modes (i.e. $\partial_{\phi_{i}}E_{\mathcal{C}}|_{\ell_{i}=\mathrm{minimizers}}=0$) and there is no energy barrier for intercalations. Consequently, there is no need for noise to drive intercalations in our model. By contrast, for the widely-used area-perimeter elastic energy $E=(A-A_{0})^{2}+(P-P_{0})^{2}$ (where $A,P$ are the cell area and perimeter, and $A_{0},P_{0}$ their target values) [Bi.etal2015], there exists an energy barrier, and the inner interface $\ell_{0}$ only collapses when $\beta\to\pi$ (see Fig. S14B). Note that the area-perimeter energy is a special case because of the geometric incompatibility of area and perimeter constraints when $P_{0}/\sqrt{A_{0}}<3.72$. For $P_{0}/\sqrt{A_{0}}>3.72$ cell shape is underdetermined (floppy) and the shear modulus vanishes. Recall that we have shown that a finite shear modulus is required to translate tension dynamics into tissue-level convergent extension (see Fig. 7F). The area-perimeter energy in the fluid regime is therefore not suitable to model convergent extension driven by internal tension dynamics.
Combining area elasticity with shear elasticity based on the shape tensor, $E=(A-A_{0})^{2}+\mu\mathrm{Tr}[(S_{\mathcal{C}}-S_{0})^{2}]$, (or perimeter elasticity with cell shape bulk elasticity) leads to similar results as (3). Note also that because of the degree-of-freedom count, the system is under-specified if the shear modulus is 0, foreshadowing the fact that without shear modulus, no convergence-extension takes place. Indeed, in the case of the symmetric lattice, the degrees of freedom are the three vertex angles, determined by the tensions, and the three interface lengths, independent of the angles. Fixing the cell area still leaves freedom to change the interface lengths. More generally, we have shown above and illustrated in Fig. 1 that isogonal modes allow for area-preserving shear transformations.
@@ 5.4 Tissue scale simulation
Using the components of the single-quartet simulation, we now describe a simulation of an arbitrary, disordered tissue patch. The key assumption is again adiabatic force balance and cortical tension dominance: the tissue geometry is found by solving for force balance, by first finding the angles from the cortical tensions and then the isogonal mode by minimizing cell elastic energy. The key quantity is the tension triangulation whose dynamics determine the tissue dynamics.
@@
We model a piece of epithelial tissue by a polygonal tiling of the plane, polygons corresponding to cells. Two cells are neighbors if they share an edge (this implies that the cell adjacency graph is a triangulation).
@@
  • Faces: primal coordinates, the $x,y$ coordinates of the vertex in the cell tessellation corresponding to the face, denoted by $\mathbf{r}_{ijk}$.
@@
The mesh is implemented in object-oriented python, using a half-edge mesh data structure [Marschner.Shirley2018]. A half-edge mesh consists of vertices, faces, and, for each edge, two oriented half-edges. A half-edge stores a reference to the next half-edge on the counter-clockwise oriented face it belongs to, and a reference to its twin, the half-edge with opposite orientation on the adjacent face. This structure defines an orientation for every triangular face and makes mesh traversal (e.g. get all vertices of a cell) and mesh modification (e.g. intercalations) very convenient.
@@
Due to the assumption of adiabatic force balance, the tissue geometry is determined by the instantaneous tensions. The dynamics are therefore determined by the change in the tension triangulation (due to positive feedback). In detail, the time evolution in our simulation is done in three steps:
@@
    Triangulation flattening: Optimize the dual vertex positions so that the length of each half-edge is as close as possible to its internally stored tension. This ensures global tension balance.
@@
    Cell shape optimization: Optimize the primal vertex positions to minimize the shape-tensor-based elastic energy, while constraining the angles by the dual triangulation (“isogonal constraint”).
@@
The next sections describe them in detail. Fig. S15 provides a summary of the simulation workflow. The most complicated and time-consuming of them is the optimization of cell shapes under the isogonal constraint. Overall, a simulation of convergent extension of $\sim 1000$ cells takes approx. 15 minutes on a current-generation laptop.
Note that the workflow of our simulations differs somewhat from the more commonly used dissipative dynamics, in which cell geometry is updated dynamically according to the ODE $\gamma\partial_{t}\mathbf{r}_{ijk}=-\frac{\partial E}{\partial\mathbf{r}_{ijk}}$. The dual separation of scales (fast relaxation and dominance of cortical tensions), means that a direct implementation of dissipative dynamics is numerically challenging. The elastic energy minimization can be seen as a relaxational Euler step with an adaptive step size.
Relaxation via fast dissipative dynamics is a viable approach that could indeed be used as an alternative numerical relaxation method. However, as the theory of fast variable elimination in the context of multiscale dynamics is well established [VanKampen1985], we do not expect a different scheme to affect our results to an important degree. The key “ingredient” of our model is not how exactly the elastic energy is relaxed, but how active tensions are regulated.
@@
The cutoff parameter $k_{\text{cutoff}}$ stops runaway positive feedback at values $T>1/k_{\text{cutoff}}$. It is not strictly necessary since runaway feedback is cut short by intercalations, but helpful for long-time simulations, where in rare cases run-away tension cable configurations occur which “blow up” numerically. We chose $1/k_{\text{cutoff}}=5\bar{T}$ where $\bar{T}$ is the overall tension scale so that it has almost no effect in normal tension configurations. We chose the feedback parameter $n=4$ and the relaxation rate $\tau_{p}=\tau_{T}/4$, the same used in the single-quartet and single-triangle simulations.
So far, the time evolution of tensions is purely autonomous (i.e. each tension triangle is independent of all others) and completely analogous to the single-quartet simulation. In the single-quartet simulation, all tension triangles are equal by symmetry and periodicity, and thus tile the plane as a triangulation by construction. In general, this is no longer guaranteed. Even if one starts with a flat tension triangulation, the autonomous time evolution (S26) will result in a set of tension triangles that cannot fit together to a flat triangulation. The triangulation will develop curvature [Noll.etal2020], i.e. “crumple” out of plane. This means that we have to add only one ingredient to enforce triangulation planarity.
We enforce planarity as follows. The set of all balanced tension triangulations is parameterized by the dual vertex positions $\mathbf{t}_{i}$. We, therefore, seek the balanced tension configuration that approximates the intrinsic tensions as closely as possible by minimizing the following elastic energy:
@@
Since the simulation is based on half-edges, the per-edge intrinsic tension $\tilde{T}_{ij}$ is the average of the tensions of the two half-edges (which can be thought of as the two cellular cortices coupled in a cell-cell interface). The second term, $\mathrm{Pen}(A,A_{0})$ is a penalty term on the triangle areas $A$ (calculated from the vertex positions $\mathbf{t}_{i}$). It fixes the overall tension scale (which is arbitrary due to force balance) and ensures the triangulation is well-behaved numerically. It contains two terms:
@@
In the simulation, we first carry out an explicit Euler step for the intrinsic tensions, and then “flatten the triangulation”, i.e. find the dual vertex positions by minimizing (S27). This determines the extrinsic tensions $T_{ij}=||\mathbf{t}_{i}-\mathbf{t}_{j}||$. Finally, we relax the intrinsic tensions to force balance with rate $\tau_{\mathrm{balance}}$,
@@
In order to ensure numerical stability even for large $\tau_{T}$, we use the exact solution of (S29) numerically, instead of carrying out an explicit Euler step.
The triangulation-flattening prescription can be thought of as modeling, indirectly, known feedback loops that lead to convergence to balanced tension, such as strain-rate feedback [Noll.etal2017], which has been demonstrated experimentally [Gustafson.etal2022]. The rate $\tau_{\mathrm{balance}}^{-1}$ determines the speed of these feedback loops. The influence of the value of $\tau_{\mathrm{balance}}^{-1}$ is studied in Fig. S9. In principle, strain rate feedback requires movement of cell vertices to sense tension imbalance and then bring tensions back to a balanced state. We neglect this vertex motion since we assume balance is restored rapidly. Directly modeling of strain rate feedback requires going beyond the separation between tension- and real-space-dynamics that is the basis of our model here and is therefore beyond the scope of this paper. Further, the triangulation-flattening prescription is the most parsimonious generalization of the local tension feedback dynamics employed in the single-quartet simulation to a disordered tissue patch.
The complete tension dynamics that are implemented in the simulation therefore comprises three terms in addition to the dynamics shown in the main text (5): (1) Numerical cutoff of positive feedback at very high tensions (2) Passive tension dynamics post T1 (3) Relaxation of tensions to force equilibrium:
@@
Above, we have described the dynamics of the tension triangulation. After an Euler step to update the tensions has been taken, we calculate the real space cell tessellation (i.e. the positions of the cell vertices) by minimizing the cell-shape elastic energy while constraining the angles by the tension triangulation. In the case of a single quartet of identical cells, we could explicitly parameterize the cell shape by interface lengths and angles, and optimize only with respect to the former. In the disordered case, it is more convenient to enforce the angle constraint via a (strong) energy penalty. It would also be possible to instead directly optimize the per-cell isogonal potentials $\Theta_{i}$ with respect to the Voronoi reference, obviating the need for a penalty.
The elastic energy to minimize therefore has two terms – the cell elastic energy (3) calculated from cell shape tensor $S_{\mathcal{C}}$ (5.3) and reference shape $S_{0}$, and the angle penalty, which corresponds to the cortical tension energy term:
@@
Here, $\hat{n}_{ij}$ is the unit normal of the tension edge, $\hat{\mathbf{r}}_{ij}$ the unit tangent vector of the primal edge, and $\overline{\ell}$ is the average interface length, to ensure the term has the correct dimensions of energy. The small parameter $\varepsilon$ (we use $\varepsilon\sim 10^{-3}\lambda^{-1}$) represents the fact that the cell shape energy is assumed to be much weaker than the cortical tensions. If desired, one can replace or modify the cell shape elastic energy $E_{\mathcal{C}}$ in (S32) by another energy, e.g. the common area-perimeter elasticity.
In (S32), we use a different term than $T_{ij}\ell_{ij}$ to represent angle constraint due to cortical tensions. When determining the vertex positions $\mathbf{r}_{ijk}$, we minimize the cell shape elastic energy under the constraint of fixed angles, as given by the tension triangulation (so in a sense, we minimize the energy not with respect to the complete vertex positions, but with respect to the isogonal modes only). The tensions are externally imposed parameters and not functions of the vertex positions, like in a spring network. (S32) does not represent the physical energy, but merely the way we numerically enforce the angle constraints. We use the alternative energy term $T_{ij}(1-\hat{n}_{ij}\cdot\hat{\mathbf{r}}_{ij})$ for two numerical reasons. (1) Compared to $T_{ij}\ell_{ij}$, it does not require a pressure term $p\sum_{i}A_{i}$ to avoid global constriction (2) it avoids potential convergence issues when minimizing the elastic energies in the presence of very short junctions, which occur before and after T1-transitions (3) it automatically penalizes invalid self-intersecting vertex configurations. However, if the boundary is fixed, both forms of the energy lead to the same family of ground states, namely those where the vertex angles are complementary to those of the tension triangulation. The energy term $T_{ij}(1-\hat{n}_{ij}\cdot\hat{\mathbf{r}}_{ij})$ was explicitly constructed to ensure this. Away from the ground state, the two energies differ; however, this is immaterial since we work in the regime where the tension constraints are always satisfied exactly. We can also rewrite
@@
After carrying out a local Euler step and updating the dual vertex positions by “triangulation flattening”, we find the primal vertex positions by minimizing the elastic energy (S32).
@@
After each time step, we check whether any edge has collapsed. We call an edge collapsed if its length is $<10\%$ of the overall length scale (mean edge length at simulation start). If this occurs, we carry out a neighbor exchange (flip the corresponding edge in the triangulation) and set the active and passive tensions on the new edge as described above.
Edges that underwent a T1 transition in the last 20 simulation time steps (corresponding to 2.5 minutes) are excluded from edge collapse to prevent the immediate re-collapse of newly formed junctions. Instead of this hard-coded criterion, one could also calculate the strain rate of junctions of short length and exclude extending junctions, as shown by the time traces of junction length in Fig. S10B.
@@
The simulation requires solving two optimization problems, one for the triangulation (S27) and one for the real-space shape (S32). To solve an optimization problem efficiently, it is crucial to know the gradient of the objective function. We use the python library JAX [Bradbury.etal2018] to automatically differentiate our energy functions, allowing easy, rapid, and bug-free calculation of the Jacobians. We further use JAX’s just-in-time functionality to accelerate our Python code. Gradient-based optimization requires the objective function to be at least once differentiable. To ensure this, we mollify functions as required; for example, the Euclidean norm $\sqrt{x^{2}+y^{2}}\rightarrow\sqrt{x^{2}+y^{2}+\varepsilon^{2}},\>\varepsilon\sim 10^{-3}$ to ensure differentiability when computing the lengths of very short edges.
@@
We consider two types of initial conditions, using either a regular lattice or a random point process to choose the tension vertices. In the former case, we chose rectangular patches of a regular triangular lattice, with i.i.d. Gaussian noise added to the vertex positions. The amplitude of the noise was chosen so that $\mathrm{std}(T_{ij})=0.1\>\mathrm{mean}(T_{ij})$. The influence of the Gaussian noise standard deviation is shown in Fig. S9. A triangular lattice corresponds to a hexagonal lattice of cells. Indeed, locally, the Drosophila embryo at the onset of GBE looks like it is composed of (noisy) hexagonal lattice patches, as shown in Fig. S1. This lattice is subjected to a small initial strain to create an initial tension anisotropy, as described in the main text.
To investigate the effect of disorder, we also create initial conditions using a hard disk process. This randomly samples 2d-points (the eventual vertices in tension space) under the constraint that no two points are closer than $2r$ where $r$ is the “hard disk” radius. Instead of $r$, we use the normalized hard disk packing fraction $\rho\in[0,1]$, which is maximal for a hexagonal lattice. At $\rho\sim 0.72$, the hard disk process undergoes a phase transition that destroys the lattice order. The limit $\rho=0$ corresponds to a Poisson point process. To sample from a hard disk point process, we use the event-chain Monte Carlo algorithm of Ref. [Bernard.etal2009], with code based on that provided with Ref. [Li.etal2022]. Note that with other algorithms such as sequential deposition or standard Monte Carlo it is either impossible or prohibitively slow to sample at arbitrary values of $\rho$.
To create a tension triangulation, we then compute the Delaunay triangulation of this point set (we further refine the initial condition by removing the edge of this tension triangulation, where the Delaunay algorithm can create very long tension edges). This triangulation is stretched uniformly to set up the initial anisotropy of tension. Cell numbers and initial aspect ratios were also chosen identically in the simulations for the phase diagram Fig. 2D.